OpenStudy (anonymous):

f(x)= sqrt(4-x) and h(x)=x^2 Find the simplified formula for this combination function and domain: (foh)(x) and (fof)(x)

OpenStudy (hba):

\[f(x)=\sqrt {4-x} \]

OpenStudy (anonymous):

I thought it was \[\sqrt{4-x^2}\]

OpenStudy (hba):

Well Doesnt Look Like That :)

OpenStudy (anonymous):

Oh okay, lol I don't get why the answer is just something thats already given lol, explainnn please? Like your steps

OpenStudy (hba):

\[(foh)(x)=\sqrt{4-(x)^2}\]

OpenStudy (hba):

\[(fof)x=\sqrt{4-(4-x)}\]

OpenStudy (hba):

Got It @jessicardoza93

OpenStudy (anonymous):

(hof)(x)?

OpenStudy (anonymous):

@hba Thanks :)

OpenStudy (hba):

\[(hof)x=(\sqrt{4-x})^2\]

OpenStudy (anonymous):

Grazie :)

OpenStudy (anonymous):

what about the domains?

OpenStudy (hba):

What Do You Mean ?

OpenStudy (anonymous):

Hey I was right on the (foh) :P hah, uhm im not sure it says "In each case say what the domain is "

OpenStudy (hba):

Please Dont Confuse Me LoL And Be Clear ;)

OpenStudy (phi):

notice that h(f(x)) simplifies to 4-x however, in the original with the square root, you must exclude x values that give you a negative number inside the root sign.

OpenStudy (anonymous):

Given: f(x)= sqrt(4-x) and h(x)=x^2 Problem: Find the simplified formula for this combination function and for each case state what the domain is a)(foh)(x) b)(fof)(x) @hba thats as clear as I can be :)

OpenStudy (anonymous):

@phi what does that mean then... ?

OpenStudy (hba):

\[(fof)x=\sqrt{4-(4-x)},(fof)x=\sqrt{-x}\]

OpenStudy (phi):

f(f(x)) is \[ \sqrt{4-\sqrt{4-x}} \] first 4-x must be ≥0 or the inside square root will be the root of a neg number 4-x≥0 4≥x now the outer root you need 4- sqrt(4-x) ≥0 4≥ sqrt(4-x) 16≥ 4-x 12≥ -x multiply by -1 and flip the ≥ -12 ≤ x so -12≤ x ≤4

OpenStudy (phi):

for (a) f(h(x)) \[\sqrt{4-x^2} \] you need to find the values of x that keep 4-x^2 ≥0

OpenStudy (anonymous):

@phi so -12≤ x ≤4 is the answer for (fof)(x)?

OpenStudy (phi):

that is the domain for (f o f) (x)

OpenStudy (anonymous):

thank you !!!